OpenStudy (anonymous):

what percent of 320 is 100

OpenStudy (anonymous):

(100/320)*100=

OpenStudy (unklerhaukus):

31.25%

OpenStudy (anonymous):

32.25

jhonyy9 (jhonyy9):

320 ...... 100 100 ...... x ---------------- x=(100*100)/320 x=1000/32 %